OpenStudy (pythagoras123):

Solve the equation 2x^3 +7x^2 -10x-24=0. Hence solve the equation 2z^3 -7z^2-10z+24 = 0.

OpenStudy (pythagoras123):

In case it was unclear: Solve the equation \[2x ^{3}+7x ^{2}-10x-24=0\] and hence solve the equation \[2z ^{3}-7z ^{2}-10z+24=0\]

OpenStudy (pythagoras123):

using a suitable substitution

OpenStudy (pythagoras123):

help @ganeshie8

OpenStudy (perl):

what kind of substitutions have you used in the past. To solve this kind of problem i would use rational roots theorem, and if that doesn't work a computer algebra system.

OpenStudy (briellshaw):

do you go to agora? @pythagoras123

OpenStudy (pythagoras123):

@BriellShaw no

OpenStudy (briellshaw):

oh nvm i see that you are in college

OpenStudy (bbmac):

Are you in college??

OpenStudy (perl):

$$2x ^{3}+7x ^{2}-10x-24=(x-2)(2x+3)(x+4)$$

OpenStudy (perl):

you can set those factors individually to zero

OpenStudy (pythagoras123):

Example of substitution would be Solve the equation \[2x ^{3}+7x ^{2}-10x-24=0\] Hence solve the equation \[2(y+2)^{3} +7(y+2)^{2}-10y-44=0\] The second equation can be written as \[2(y+2)^{3}+7(y+2)^{2}-10(y+2)-24=0\] Hence substituting x as (y+2)

OpenStudy (pythagoras123):

Thing is, I don't know how to do the substitution for the specific equation I listed

OpenStudy (perl):

There are substitutions used to solve cubic equations.

OpenStudy (perl):

I was looking at this http://home.scarlet.be/math/cubic.htm